# Dimension-reduced Reconstruction Map Learning for Parameter Estimation in Likelihood-Free Inference Problems

### v1

@@ 3.4 Evaluating Estimation Performance
$R(\theta,\widehat{\theta}):=E^{y\mid\theta}(l(\theta,\widehat{\theta}(y))),$ (8)
@@
$r(\rho,\widehat{\theta}):=E^{(y,\theta)}(l(\theta,\widehat{\theta}(y)))=E^{\theta\sim\rho}(R(\theta,\widehat{\theta})),$ (9)
@@ 4 Numerical Experiments
We consider four simulated examples to evaluate the performance of the proposed estimation framework. The first three examples feature an intractable likelihood, while the final example is defined by a highly nonlinear and nonconvex likelihood surface, which poses computational challenges to likelihood-based methods. An accessible likelihood allows us to compare RM and RM-DR’s performance with maximum likelihood estimation and to demonstrate the RM-DRLO approach in the supplement. The RM, RM-DR, and RM-DRLO approaches are based on a fully-connected neural network with 2 hidden layers, each having 32 neurons, and a ReLU activation function. The choice of this NN architecture involves trial and error through evaluating performance on synthetic validation data, and other more complex NN architectures produce comparable performance (which we do not show here for brevity). The design distribution for parameters is taken to be a uniform distribution over the parameter space. The training sample consists of 125,000 output-parameter pairs, of which $25\%$ is held out for validation. Evaluation is based on the fit criteria discussed in Section 3.4, which are approximated based on $L=100$ replications. ABC is implemented using the parallel tempering algorithm for efficient posterior exploration (Swendsen and Wang, 1986; Geyer, 1991). A Gaussian kernel is used to measure the similarity between observed and synthetic data, with the bandwidth parameter chosen manually to be as small as possible while resulting in an acceptance rate within the target range. Convergence is assessed by monitoring traceplots and correlation plots. Finally, MLE and SLE estimators are obtained via numerical optimization using the dual annealing algorithm.
@@ 4.1 Ricker Model
where $\delta$ is an unknown scale parameter. The initial population is $N(0)=2$, and data $y=(y(1),\ldots,y(1,000))^{\top}$ is observed at $m=1,000$ consecutive time steps. Setting $\eta=\log(a)$, the parameters of interest are $\theta=(\eta,\sigma,\delta)^{\top}\in(2,5)\times(0,0.3)\times(1,4)$. Supplement Fig. 1 shows four replications of $y$ under $\theta=(3,0.2,2)^{\top}$, illustrating the diversity of sample paths that are possible under the same parameter setting. A likelihood calculation would require marginalization over $m$ unobserved population densities, and is thus effectively intractable.
@@
RM-DR, SLE, and ABC are implemented using the summary statistics suggested in Wood (2010). Let $\Delta(t)=y(t)-y(t-1)$ denote the differences between consecutive observations and let ${\Delta}_{(t)}$ be the $t$-th order statistics of $\Delta(1),\dots,\Delta(1000)$. Similarly, let $y_{(t)}$ be $t$-th order statistics of $y(1),\dots,y(1,000)$. The summary statistics are: the sample mean $\overline{y}=\frac{1}{1,000}\sum_{t=1}^{1000}y(t)$, sample autocovariance $\upsilon(h)=\frac{1}{1,000}\sum_{t=1}^{1,000-h}(y(t+h)-\overline{y})(y(t)-\overline{y})$ with lag $h$ from 00 to $5$, number of zeros observed $\tau=\sum_{t=1}^{1,000}1(y(t)=0)$, coefficients of the cubic regression of ordered differences $\Delta^{(t)}$ on the ordered observed values $y^{(t)}$, and coefficients of the autoregression of ${(y(t+1))}^{0.3}$ on ${y(t)}^{0.3}$ and ${y(t)}^{0.6}$. The rationale for these choices is as follows. The sample mean and autocovariance are typically useful summaries of time series data, while the frequency of zero observations can provide insights into the distribution of Poisson data. Coefficients of the cubic regression can summarize the marginal distribution of observations, and coefficients of the autoregression contain information about dynamic structure.
@@
Fig. 2 shows scatter plots of estimates versus simulation values of $\eta,\sigma$ and $\delta$ (rows), respectively, under the four different estimation methods (columns). Out of the approaches considered, RM-DR estimates are the closest to true values for all components of $\theta$ (smallest spread around the 45° line, in red). RM achieves the worst performance among all methods considered. ABC shows comparable estimation accuracy to SLE, except for the estimation of the standard deviation $\sigma$. Supplement Fig. 2 compares the performance of RM and RM-DR estimators using the evaluation criteria introduced in Section 3.4. Each point on the 3-d plots corresponds to one of 1,000 different simulation parameter setting. The color corresponds to the magnitude of the log squared bias, variance, and MSE (rows), respectively, for RM (left column) and RM-DR (right column). RM-DR estimators achieve lower squared bias, variance and MSE across almost all parameter values compared to RM. Indeed, RM-DR achieves substantially lower integrated squared bias, variance, and MSE. Due to the relatively expensive computation (approximately several hours for a single estimate), the performance of ABC and SLE are compared under three different $\theta$ settings in Table 1. RM-DR has the lowest MSE, variance, and squared bias in almost all cases, followed by ABC and SLE. RM has the worst performance over all metrics. In summary, this example illustrates that summarization of the data informing the RM-DR method greatly improves estimation performance relative to the RM approach, and performs favorably relative to ABC and SLE both in terms of accuracy and speed under the same choice of summaries.
@@ 4.2 M/G/1-queue
We next consider a queuing model consisting of a first-come-first-serve single-server queue (M/G/1-queue), used in Fearnhead and Prangle (2012) as an example of a stochastic simulation model with an intractable likelihood. The service times are uniformly distributed on the interval $\left[\theta_{1},\theta_{2}\right]$, and inter-arrival times are exponentially distributed with rate $\theta_{3}$. The simulation procedure for the $n$th inter-departure time $y(n)$ is provided in the supplement. Assume that the first 1,000 inter-departure times $y=(y(1),\ldots,y(1,000))^{\top}$ are observed and the parameters $\theta=(\theta_{1},\theta_{2},\theta_{3})^{\top}$ are unknown. Supplement Fig. 3 shows a histogram of the inter-departure times from four independent realizations (panels) of $y$ when $\theta=\left(4,8,1/6\right)^{\top}$. The design distribution for $(\theta_{1},\theta_{2}-\theta_{1},\theta_{3})^{\top}$ is chosen as a uniform distribution over the region $\left(0,10\right)\times\left(0,10\right)\times\left(0,\frac{1}{3}\right)$, where the resulting service and inter-arrival times have on average comparable magnitudes.
@@
Fig. 3 shows scatter plots of estimates versus simulation values of the components of $\theta$ (rows), under the four different estimation methods (columns). The smallest spread of values around the 45° line (red) indicating correct estimation is achieved by RM-DR. As in the previous example, supplement Fig. 4 shows a comparison between the performance of RM and RM-DR estimators. On average, RM-DR achieves substantially better estimation performance than RM, in terms of squared bias, variance, and MSE across all the simulation parameter values. As in the previous example, SLE and ABC are further evaluated at three $\theta$ settings in Table 2. RM-DR has the best performance across all criteria, followed by ABC. The estimation performance of RM and SLE is substantially worse, with much higher values of squared bias and variance. Looking at MSE, RM performs marginally better than SLE, mainly due to lower estimation variance. Once again, the RM-DR method performs well in the likelihood-free setting under a sensible choice of summary statistics.
@@ 4.3 Lotka–Volterra Model
where $u$ and $v$ represent the abundance of a prey and predator species, respectively. The first reaction describes prey production (e.g., through birth or immigration), the second reaction captures consumption of prey by the predator, and the third reaction represents removal of predators (e.g. through death or out-migration). These dynamics can be described by a continuous-time discrete state Markov chain, where each reaction occurs at a rate that depends on the current state of the system, specified in terms of transition probabilities over a small time interval $\left(t,t+\delta t\right]$, as explained in the supplement. We denote the state of the system at time $t$ by $y(t)=\left(u(t),v(t)\right)^{\top}$, where $u(t)$ and $v(t)$ represent the abundance of prey and predators at time $t$, respectively. Assume the initial condition $y(0)=(50,100)^{\top}$ and unknown parameters $\theta=(\theta_{1},\theta_{2},\theta_{3})^{\top}\in(0.3,0.6)\times(0.005,0.01)\times(0.1,0.4)$. In this example, we observe both prey and predator populations at 1,000 equidistant points in a time interval $\left[0,30\right]$, and denote the observed abundances by $u\in\mathbb{R}^{1,000}$ and $v\in\mathbb{R}^{1,000}$, respectively. Given $\theta$, we simulate data using the Gillespie algorithm (Gillespie, 1977), as illustrated in supplement Fig. 5.
@@
Looking at the scatter plots of estimated values versus simulation values of parameters in Fig. 4, RM-DR estimates are both more accurate and less variable than RM and SLE. While ABC has low variability overall, it does have high bias. Performance of RM and RM-DR is illustrated in supplement Fig. 6 at 200 uniformly sampled $\theta$ settings. RM-DR estimation has lower squared bias, variance and MSE for most $\theta$ values relative to RM estimation. It also has lower integrated versions of these metrics, which demonstrates a better overall performance. Again we evaluate performance of ABC and SLE at several $\theta$ settings in Table 3, showing that RM-DR has the lowest MSE and squared bias across the estimation methods considered. ABC is second to RM-DR in terms of MSE, mainly due to having low variance, while SLE has the worst overall performance.
@@ 4.4 FitzHugh–Nagumo Model
The final numerical example considers parameter estimation for the FitzHugh–Nagumo (FN) ODE model, which describes the time evolution of voltage $v(t)$ and recovery $r(t)$ across the membrane of a biological neuron. The ODE initial value problem (see supplement) depends on unknown parameters $\theta=(\theta_{1},\theta_{2})^{\top}$, fixed constants $\tau=3$ and $\zeta=0.4$, and initial conditions $v(0)=r(0)=0$. We make the standard assumption that only voltage is observed with additive noise via $y(t)=v(t)+\epsilon$ at a discrete set of locations $t_{i}=0.025i$ for $i=1,\ldots,1,000$, and with $\epsilon\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\mathcal{N}(0,0.06^{2})$. The ODE solution $v(t)$ under different parameter settings is shown in supplement Fig. 7. Since the likelihood for this model can be approximated numerically, we can compare RM and RM-DR estimation with maximum likelihood estimation. The summaries chosen for RM-DR implementation are the coefficients of a nonlinear regression on $K$ Fourier basis functions, chosen as a way of extracting frequency and amplitude information from this periodic system.
@@
For the integrated performance metrics, we consider test parameter values over the grid $(\theta_{1},\theta_{2})\in\{-0.2+0.03j\,,\,-0.4+0.04l\}_{j,l=0,1,\ldots,40}$. We vary the number of basis functions $K$ to investigate how the summary dimension impacts RM-DR performance. Fig. 5 shows that RM-DR has the smallest integrated squared bias, variance, and MSE of the three methods, regardless of input dimension. Its performance is robust to different choices of input dimension within a reasonable range for this example. As expected, using $K=5$ basis coefficients leads to notably worse performance than using larger values of $K$, as the latter choices convey more amplitude and frequency information. Unsurprisingly, RM has the largest integrated squared bias and variance due to the difficulty in reconstructing a mapping with a large input space. Supplement Fig. 8 shows Monte Carlo estimates of log squared bias, variance, and MSE for the three methods considered. For all approaches, estimation is worse for simulation parameter values in the top left triangular region of the parameter space. This is because the ODE solution associated with these parameter quickly attains a steady state, as shown in supplement Fig. 7, containing less information about the parameter. RM-DR provides an improvement over RM at many parameter values. Comparing RM-DR with MLE, RM-DR has better estimations in the top left region of the parameter space. Although MLE produces lower bias and better estimation at most parameter values, it has larger variance in the top left region, which leads to a higher IMSE relative to RM-DR. Because the likelihood is available in this example, we can also test the RM-DRLO approach described in the supplement, which consists of using RM-DR as a starting for to a local optimization algorithm as an alternative to an expensive global optimization method. The results are described in supplement section 4.4.
@@ 5 Summary
We propose a simulation-based RM-DR estimation approach with dimension reduction for the class of inverse problems in which a closed-form likelihood is unavailable or expensive, and discuss its properties, evaluation criteria, and uncertainty quantification. This approach resolves the problem of degraded estimation performance when data dimension increases, which makes direct reconstruction map estimation unreliable in practice. We show that under certain assumptions, the RM-DR estimator is asymptotically equivalent to a Bayes estimator. By learning a dimension-reduced manifold, RM-DR reduces the approximation error relative to RM estimation, as illustrated in multiple numerical experiments. Additionally, in the setting where the likelihood is available but expensive, we propose to combine the RM-DR approach with local optimization methods as an alternative to global optimization approaches for parameter estimation, with comparable accuracy but more time efficiency.
@@
Future work includes investigating large-sample properties of the RM-DR estimator, its performance as both the training sample size and the dimension of summary statistics grows, and ways to choose appropriate summary statistics.

### v2

@@ 3.4 Evaluating Estimation Performance
$R(\theta,\widehat{\theta}):=\mathbb{E}_{y\mid\theta}\bigl{[}l(\theta,\widehat{\theta}(y))\bigr{]},$ (9)
@@
$r(\rho,\widehat{\theta}):=\mathbb{E}_{(y,\theta)}\bigl{[}l(\theta,\widehat{\theta}(y))\bigr{]}=\mathbb{E}_{\theta\sim\rho}\bigl{[}R(\theta,\widehat{\theta})\bigr{]},$ (10)
@@ 4 Numerical Experiments
[…]
@@ 4.1 Ricker Model
where $\delta$ is an unknown scale parameter. The initial population is $N(0)=2$, and data $y=(y(1),\ldots,y(1,000))^{\top}$ is observed at $m=1,000$ consecutive time steps. Setting $\eta=\log(a)$, the parameters of interest are $\theta=(\eta,\sigma,\delta)^{\top}\in(2,5)\times(0,0.3)\times(1,4)$. Supplement Fig. S1 shows four replications of $y$ under $\theta=(3,0.2,2)^{\top}$, illustrating the diversity of sample paths that are possible under the same parameter setting. A likelihood calculation would require marginalization over $m$ unobserved population densities, and is thus effectively intractable.
@@
RM-DR, SLE, and ABC are implemented using the summary statistics suggested in Wood (2010). Let $\Delta(t)=y(t)-y(t-1)$ denote the differences between consecutive observations and let ${\Delta}_{(t)}$ be the $t$-th order statistics of $\Delta(1),\dots,\Delta(1000)$. Similarly, let $y_{(t)}$ be $t$-th order statistics of $y(1),\dots,y(1,000)$. The summary statistics are: the sample mean $\overline{y}=\frac{1}{1,000}\sum_{t=1}^{1000}y(t)$, sample autocovariance $\upsilon(h)=\frac{1}{1,000}\sum_{t=1}^{1,000-h}(y(t+h)-\overline{y})(y(t)-\overline{y})$ with lag $h$ from 0 to 555, number of zeros observed $\tau=\sum_{t=1}^{1,000}1(y(t)=0)$, coefficients of the cubic regression of ordered differences $\Delta^{(t)}$ on the ordered observed values $y^{(t)}$, and coefficients of the autoregression of ${(y(t+1))}^{0.3}$ on ${y(t)}^{0.3}$ and ${y(t)}^{0.6}$. The rationale for these choices is as follows. The sample mean and autocovariance are typically useful summaries of time series data, while the frequency of zero observations can provide insights into the distribution of Poisson data. Coefficients of the cubic regression can summarize the marginal distribution of observations, and coefficients of the autoregression contain information about dynamic structure.
@@
Fig. 2 shows scatter plots of estimates versus simulation values of $\eta,\sigma$ and $\delta$ (rows), respectively, under the four different estimation methods (columns). Out of the approaches considered, RM-DR estimates are the closest to true values for all components of $\theta$ (smallest spread around the 45° line, in red). RM achieves the worst performance among all methods considered. ABC shows comparable estimation accuracy to SLE, except for the estimation of the standard deviation $\sigma$. Supplement Fig. S2 compares the performance of RM and RM-DR estimators using the evaluation criteria introduced in Section 3.4. Each point on the 3-d plots corresponds to one of 1,000 different simulation parameter setting. The color corresponds to the magnitude of the log squared bias, variance, and MSE (rows), respectively, for RM (left column) and RM-DR (right column). RM-DR estimators achieve lower squared bias, variance and MSE across almost all parameter values compared to RM. Indeed, RM-DR achieves substantially lower integrated squared bias, variance, and MSE. Due to the relatively expensive computation (approximately several hours for a single estimate), the performance of ABC and SLE are compared under three different $\theta$ settings in Table 1. RM-DR has the lowest MSE, variance, and squared bias in almost all cases, followed by ABC and SLE. RM has the worst performance over all metrics. In summary, this example illustrates that summarization of the data informing the RM-DR method greatly improves estimation performance relative to the RM approach, and performs favorably relative to ABC and SLE both in terms of accuracy and speed under the same choice of summaries.
@@ 4.2 M/G/1-queue
We next consider a queuing model consisting of a first-come-first-serve single-server queue (M/G/1-queue), used in Fearnhead and Prangle (2012) as an example of a stochastic simulation model with an intractable likelihood. The service times are uniformly distributed on the interval $\left[\theta_{1},\theta_{2}\right]$, and inter-arrival times are exponentially distributed with rate $\theta_{3}$. The simulation procedure for the $n$th inter-departure time $y(n)$ is provided in the supplement. Assume that the first 1,000 inter-departure times $y=(y(1),\ldots,y(1,000))^{\top}$ are observed and the parameters $\theta=(\theta_{1},\theta_{2},\theta_{3})^{\top}$ are unknown. Supplement Fig. S3 shows a histogram of the inter-departure times from four independent realizations (panels) of $y$ when $\theta=\left(4,8,1/6\right)^{\top}$. The design distribution for $(\theta_{1},\theta_{2}-\theta_{1},\theta_{3})^{\top}$ is chosen as a uniform distribution over the region $\left(0,10\right)\times\left(0,10\right)\times\left(0,\frac{1}{3}\right)$, where the resulting service and inter-arrival times have on average comparable magnitudes.
@@
Fig. 3 shows scatter plots of estimates versus simulation values of the components of $\theta$ (rows), under the four different estimation methods (columns). The smallest spread of values around the 45° line (red) indicating correct estimation is achieved by RM-DR. As in the previous example, supplement Fig. S4 shows a comparison between the performance of RM and RM-DR estimators. On average, RM-DR achieves substantially better estimation performance than RM, in terms of squared bias, variance, and MSE across all the simulation parameter values. As in the previous example, SLE and ABC are further evaluated at three $\theta$ settings in Table 2. RM-DR has the best performance across all criteria, followed by ABC. The estimation performance of RM and SLE is substantially worse, with much higher values of squared bias and variance. Looking at MSE, RM performs marginally better than SLE, mainly due to lower estimation variance. Once again, the RM-DR method performs well in the likelihood-free setting under a sensible choice of summary statistics.
@@ 4.3 Lotka–Volterra Model
[…]
@@
Looking at the scatter plots of estimated values versus simulation values of parameters in Fig. 4, RM-DR estimates are both more accurate and less variable than RM and SLE. While ABC has low variability overall, it does have high bias. Performance of RM and RM-DR is illustrated in supplement Fig. S6 at 200 uniformly sampled $\theta$ settings. RM-DR estimation has lower squared bias, variance and MSE for most $\theta$ values relative to RM estimation. It also has lower integrated versions of these metrics, which demonstrates a better overall performance. Again we evaluate performance of ABC and SLE at several $\theta$ settings in Table 3, showing that RM-DR has the lowest MSE and squared bias across the estimation methods considered. ABC is second to RM-DR in terms of MSE, mainly due to having low variance, while SLE has the worst overall performance.
@@ 4.4 FitzHugh–Nagumo Model
The final numerical example considers parameter estimation for the FitzHugh–Nagumo (FN) ODE model, which describes the time evolution of voltage $v(t)$ and recovery $r(t)$ across the membrane of a biological neuron. The ODE initial value problem (see supplement) depends on unknown parameters $\theta=(\theta_{1},\theta_{2})^{\top}$, fixed constants $\tau=3$ and $\zeta=0.4$, and initial conditions $v(0)=r(0)=0$. We make the standard assumption that only voltage is observed with additive noise via $y(t)=v(t)+\epsilon$ at a discrete set of locations $t_{i}=0.025i$ for $i=1,\ldots,1,000$, and with $\epsilon\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\mathcal{N}(0,0.06^{2})$. The ODE solution $v(t)$ under different parameter settings is shown in supplement Fig. S7. Since the likelihood for this model can be approximated numerically, we can compare RM and RM-DR estimation with maximum likelihood estimation. The summaries chosen for RM-DR implementation are the coefficients of a nonlinear regression on $K$ Fourier basis functions, chosen as a way of extracting frequency and amplitude information from this periodic system.
@@
For the integrated performance metrics, we consider test parameter values over the grid $(\theta_{1},\theta_{2})\in\{-0.2+0.03j\,,\,-0.4+0.04l\}_{j,l=0,1,\ldots,40}$. We vary the number of basis functions $K$ to investigate how the summary dimension impacts RM-DR performance. Fig. 5 shows that RM-DR has the smallest integrated squared bias, variance, and MSE of the three methods, regardless of input dimension. Its performance is robust to different choices of input dimension within a reasonable range for this example. As expected, using $K=5$ basis coefficients leads to notably worse performance than using larger values of $K$, as the latter choices convey more amplitude and frequency information. Unsurprisingly, RM has the largest integrated squared bias and variance due to the difficulty in reconstructing a mapping with a large input space. Supplement Fig. S8 shows Monte Carlo estimates of log squared bias, variance, and MSE for the three methods considered. For all approaches, estimation is worse for simulation parameter values in the top left triangular region of the parameter space. This is because the ODE solution associated with these parameter quickly attains a steady state, as shown in supplement Fig. S7, containing less information about the parameter. RM-DR provides an improvement over RM at many parameter values. Comparing RM-DR with MLE, RM-DR has better estimations in the top left region of the parameter space. Although MLE produces lower bias and better estimation at most parameter values, it has larger variance in the top left region, which leads to a higher IMSE relative to RM-DR. Because the likelihood is available in this example, we can also test the RM-DRLO approach described in the supplement, which consists of using RM-DR as a starting for to a local optimization algorithm as an alternative to an expensive global optimization method. The results are described in supplement section 4.4.
@@ 5 Summary
We propose a simulation-based RM-DR estimation approach with dimension reduction for the class of inverse problems in which a closed-form likelihood is unavailable or expensive, and discuss its properties, evaluation criteria, and uncertainty quantification. This approach resolves the problem of degraded estimation performance when data dimension increases, which makes direct reconstruction map estimation unreliable in practice. We show that under mild assumptions, the RM-DR estimator converges to a Bayes estimator in probability. By learning a dimension-reduced manifold, RM-DR reduces the approximation error relative to RM estimation, as illustrated in multiple numerical experiments. Additionally, in the setting where the likelihood is available but expensive, we propose to combine the RM-DR approach with local optimization methods as an alternative to global optimization approaches for parameter estimation, with comparable accuracy but more time efficiency.
@@
In terms of future work, for high-dimensional data where informative application-specific summaries are not readily available—particularly in complex or less interpretable settings—we propose exploring the automatic learning of summary statistics. This could involve using unsupervised learning techniques, such as autoencoders, to reduce the dimensionality of the data before inputting it into the neural network model. Alternatively, we may incorporate a transformer encoder directly into the neural network architecture to summarize key features in sequential data. By leveraging the transformer’s self-attention mechanism to capture long-range dependencies and contextual relationships, the model may learn richer representations, potentially improving the robustness and accuracy of parameter estimation.
Data Availability Statement
All numerical experiments were based on a large number of simulated datasets. No real data was analyzed as part of this work.